OpenStudy (anonymous):

The parabola with the equation y = x2 was both horizontally and vertically translated to produce the parabola with the equation y = x2 – 14x + 39. Which statement is true? The parabola with the equation y = x2 was translated 7 units to the left and 10 units down. B) The parabola with the equation y = x2 was translated 7 units to the left and 10 units up. C) The parabola with the equation y = x2 was translated 7 units to the right and 10 units down. D) The parabola with the equation y = x2 was translated 7 units to the right and 10 units u

OpenStudy (anonymous):

y=x2-14x+39 standard equation y = ax^2 after shifting to (x1,y1) new equation y-y1=a*(x-x1)^2 convert your equation in this form

OpenStudy (anonymous):

y=x2-2*7*x+39+10-10 10 is added and subtracted to make it perfect square y=(x2-2*7*x+49)-10 y+10=(x2-2*7*x+7^2) y-(-10)=(x-7)^2 so shifted to (7,-10)

OpenStudy (anonymous):

so 7 unit right and 10 unit down